# Akaki Mamageishvili Offchain Labs Benny Sudakov ETH Zurich

### v2

@@ Abstract
We compare the efficiency of secure restaking and Proof-of-Stake (PoS) protocols in terms of stake requirements. First, we consider the sufficient condition for the restaking graph to be secure. We show that the condition implies that it is always possible to transform such a restaking graph into separate secure PoS protocols. Next, we derive two main results, giving upper and lower bounds on required extra stakes that one needs to add to validators of the secure restaking graph to be able to transform it into secure PoS protocols. In particular, we show that the restaking savings compared to PoS protocols can be very large and can asymptotically grow in the worst case as a square root of the number of validators. We also study a complementary question of aggregating secure PoS protocols into a secure restaking graph and provide lower and upper bounds on the PoS savings compared to restaking.
@@ 1 Introduction
Economic security of blockchains evolves over time. Initially, Proof-of-Work was proposed for Bitcoin security, [6]. Soon after that, Proof-of-Stake was proposed as an energy-efficient alternative to the Proof-of-Work concept  [^1] See [7] for a more detailed discussion on Proof-of-Work vs. Proof-of-Stake concepts.. More recently, Restaking has been proposed as a cryptoeconomically more efficient alternative to the Proof-of-Stake solution concept for blockchains. Restaking refers to a setting in which validators that are staked for one protocol can reuse their stake to secure other projects or protocols of their choice. This creates a bipartite graph structure between the projects and the validators.
@@
It can be argued that restaking gives a boost to newer projects in attracting capital for their cryptoeconomic security. Cryptoeconomic security refers to a security derived from staked parties being slashed their stakes in case of misbehavior. Restaking seems to be attractive for the validators of the original protocol as well. In particular, restaking allows them to reuse their computational power and storage allocation for additional payoffs coming from other projects’ rewards. The original protocol, for which the validators are staked, may have requirements that do not exhaust all the resources of the validators, usually for the purpose of decentralization [^2] Check out Ethereum network plans to resolve this asymmetry between validator resources using rainbow staking https://ethresear.ch/t/unbundling-staking-towards-rainbow-staking/18683.. Another potential reason for such availability might be economies of scale for establishing further validation services.
@@
The simplest way to verify the security of the restaking uses a trivial cost-benefit analysis. In this analysis, the total amount of lost stakes of attackers is compared to the total value obtained from the attack of a set of validators. The same cost-benefit analysis can be used to check the security of a Proof-of-Stake protocol, in which a single project/service is secured by multiple validators. This allows a reasonable comparison of restaking and multiple PoS protocols. [2] discusses ways to estimate the value gained from attacking a protocol and cryptoeconomic security more in detail.
@@
Since the Proof-of-Stake protocol is simpler and hence a more natural solution concept than restaking, we are interested in checking comparative benefits of restaking protocols. One particular attractive point of restaking protocols can be the lower total staking requirements than that of Proof-of-Stake protocols, since there are opportunity costs associated with locking up stake for security reasons.
@@
Most PoS protocols pay for this cost by rewarding stakers with newly issued tokens, which leads to inflation. Hence, minimizing the total stake amount while maintaining security can be a reasonable goal in both restaking or PoS protocols. Starting with a secure restaking protocol and initial stake endowments for all validators, we measure how much extra stake would be required if validators decided to form separate Proof-of-Stake protocols with their stake amounts, dividing them across the same set of projects they secure in the original restaking graph.
@@
The security parameter of a service establishes what share of the total stake securing the service is needed to successfully attack this service. Let us first check that the restaking graph is secure. It is easy to see:
@@
  • the second validator alone can not attack any service: it is not securing th first service, while the second service has large enough security parameter.
  • and, finally, both validators can attack both services but they lose the same amount of stake as they gain the value.
@@
Figure 1: Example construction for $m=3$.
We only allow all validator stakes to (weakly) increase. This captures a real-life consideration in which validators’ stakes can not be taken away, but their stakes can be increased through financing when needed. If both increasing and decreasing of validator stakes were allowed, it would be trivial to construct a restaking graph with minimum total stake. This amount is equal to the sum of all project values, and secure PoS protocols for all projects can be constructed by dividing the stakes of the validators.
@@
    We check the sufficient condition for a secure restaking graph, specified by the EigenLayer project [8]. This sufficient condition is the only currently available (generic) condition that can be checked in polynomial time in the size of the restaking graph. Restaking graphs that satisfy the sufficient condition do not save on the total required stake. In particular, it is possible to divide the initial stakes of the validators to obtain secure PoS protocols across all projects.
@@
    To further extend the analysis, we define restaking savings of a secure restaking graph. It is equal to the extra stake required, relative to the original total available stake in the graph, to be able to obtain secure PoS protocols through the division of stake described above. Higher restaking savings indicate an advantage of restaking to PoS.
@@
    We study a complementary question of aggregating given secure PoS protocols into a secure restaking graph. When aggregating, each validator adds stakes across projects that they secure in separate PoS protocols, resulting in a restaking graph. Similarly to the restaking savings, we define PoS savings as the total additional stake required to add validators in the aggregate restaking graph to make it secure. We first construct an example showing that the resulting restaking graph is not secure, in fact, at least an additional stake equal to the original total stake is required to make it secure, thus showing a lower bound of $1$ on PoS savings. We also provide an upper bound on PoS savings as a function of the underlying restaking graph.
@@
The idea of restaking was originally proposed in the whitepaper of the Eigenlayer project, [8]. A similar concept existed in traditional finance literature under the name of rehypothecation[4]. The interdependencies between protocols and validators introduced by restaking create new risks, which have been modeled and analyzed in the original whitepaper, as well as in a follow up paper [3], in which the authors study robust restaking networks, resistant to a failure of a fraction of validating stakes.  [1] studies a setting in which the value of attacked set of services is not the sum of values in the set, but a concave function. [9] studies elastic restaking services, in which allocated stakes to services make up more than the total stake. [5] studies risks and benefits of combining different restaking services.
@@ 2 Notation and Model
Our notation closely follows [3]. A restaking graph $G=(S,\pi,\alpha,V,\sigma,E)$ consists of the following:
  • $S$ denotes the set of $m$ services $\{1,...,m\}$,
  • $\pi\in\mathbb{R}^{m}_{+}$ denotes values of services (we will refer projects as services from now on throughout the paper),
@@
Let $N_{G}(s)$ denote the neighborhood of the service $s\in S$ in the restaking graph $G$, i.e. $N_{G}(s):=\{v\in V:(s,v)\in E\}$, is the set of validators staked on service $s$. Similarly, $N_{G}(v)$ denotes the neighborhood of the validator $v\in V$, i.e., $N_{G}(v):=\{s\in S:(s,v)\in E\}$, is the set of services validator $v$ is staked on.
@@
A division of stake $\sigma_{v}$ of validator $v$ across services it secures in $G$ induces a vector $c^{v}\in\mathbb{R}^{|N_{G}(v)|}_{+}$, such that, $\sum_{s\in N_{G}(v)}c^{v}_{s}=\sigma_{v}$. We say that division of stake is secure if in the corresponding Proof-of-Stake protocols, that represent star graphs centered at the service, we have guaranteed security. That is, for each service $s$ with security parameter $\alpha_{s}$, a vector of stakes $\{k\in N_{G}(s):c^{k}_{s}\}$ is secure, which means that there is no profitable attack by validators having stakes $\{k\in N_{G}(s):c^{k}_{s}\}$. Let $\mathcal{W}(G)$ denote all initial stake vectors $\sigma$ for which such division is possible. To simplify notation, here we implicitly assume that the stake vector is not part of a restaking graph $G$, and only keep its combinatorial nature together with the $\pi$ and $\alpha$ vectors.
@@ 3 Results
In this section, we first check the sufficient condition (1), then define restaking and Proof-of-Stake savings, and last, we obtain lower- and upper-bound results for them.
@@ 3.1 Restaking Savings
stake to a service $s$ it secures in $G$. Such allocation is possible because of (1). Each service $s$ is assigned a total stake amount of at least $\frac{\pi_{s}}{\alpha_{s}}$, since
@@
Note that the value $\frac{\pi_{s}}{\alpha_{s}}$ is the minimum required total stake in a PoS protocol for a service $s$ with value $\pi_{s}$ and security parameter $\alpha_{s}$ that guarantees that there is no profitable attack. For any smaller amount $t<\frac{\pi_{s}}{\alpha_{s}}$, there exists a distribution of the total stake $t$ among any number of validators such that there is a profitable attack.
There is a minimum total stake requirement to ensure that there is a division in secure PoS protocols.
@@
Another consideration is that the restaking graph with the minimum total stake in ${\cal W}(G)$ is not necessarily secure, while we will focus on secure restaking graphs from now on.
Next, we define the restaking savings for a given secure restaking graph:
@@
where $\mathbb{I}(P)=1$ whenever the statement $P$ is true.
@@
    the vector of stakes $(c_{1},...,c_{d_{G}(s)})$ is equal to $(0,...,0,\pi_{s}+\sigma_{v},0,...0)$ for any $v$. This construction is secure for a similar reason to above.
@@
Next, we show $RS(G)\leq K$. Consider a cover of vertices in $V$ by neighborhoods of services $s\in R\subseteq S$ that has the minimum maximum incidence number $K$. For each neighborhood $N_{G}(s)$, where $s\in R$, use construction 3: original $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ stakes from the restaking graph to secure service $s$ in the PoS protocol. To be able to do that, we need to add at most K1K-1 times all $\sigma_{v}$ stakes, for any $v\in V$. In this way, we utilize all $\sigma_{v}$ stakes and secure services in the cover set $R$. For service $s$ in $S\setminus R$, we add $\pi_{s}$ stake once to any of its neighbors in $N_{G}(s)$ and use construction $(0,,,0,\pi_{s},0,...,0)$ to secure service $s$ in a corresponding PoS protocol. This makes sure we have secured all services and utilized all stakes of all validators. Then, $RS(G)\leq(K-1)+1=K$. ∎
Note that $RS(G)\leq K$ in particular implies that $RS(G)\leq\max_{s\in S}(d_{G}(s))$, as $K$ is upper bounded by $\max_{s\in S}(d_{G}(s))$. There are examples where $K=\max_{s\in S}(d_{G}(s))$, implying that the upper bound can be as high as $n$. However, next, we show an upper bound as a function of the number of validators, which is asymptotically much lower than $n$.
@@
The proof combines constructions from the proof of Proposition 3. Initialize the set $S_{L}:=\emptyset$ and $V_{L}:=\emptyset$. The following procedure is repeated until it is no longer possible: At each step, find a service $s$ in $S\setminus S_{L}$ such that its degree in the set $V\setminus V_{L}$ is at least $\sqrt{n}$. Update the set $S_{L}:=S_{L}\cup s$ and $V_{L}:=V_{L}\cup N_{G}(s)$. Note that after at most $\sqrt{n}$ steps, it is not possible to find a new service $s$ and, hence, the procedure stops. That is, for the resulting set of services $S_{L}$, we have $|S_{L}|\leq\sqrt{n}$. This, in particular, implies that the maximum outdegree of a validator $v$ in $V_{L}$ to $S_{L}$, denoted by $d_{S_{L}}(v)$, is upper bounded by $\sqrt{n}$.
@@
Note that in the proofs of Proposition 3 and Theorem 1, we did not use any property of $\alpha$ values. Both upper bounds use combinatorial features of the underlying restaking graph $G$. Next, we show an upper bound that is a function of the security parameters $\alpha_{s}$. We show the following upper bound on $RS(G)$:
@@
Assume that services are labeled in the way that the security parameters are sorted in increasing order $\alpha_{1}\leq\alpha_{2}\leq...\leq\alpha_{m}$. Then, the following inequality $RS(G)\leq\frac{1}{\alpha_{2}}$ holds.
@@
For the service $s=1$, allocate the original stakes of the restaking graph in the PoS protocol: $(c_{1},...,c_{d_{G}(1)})=(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(1)}})$. This PoS protocol is secure. Repeat the following procedure for all services $s>1$:
@@
    use construction $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ to secure service $s$ in PoS protocol,
@@
    set all values in the vector $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ to 0.
Each PoS derived at step 2. is secure, as the total stake allocated for a service $s>1$ is at least $\frac{\pi_{s}}{\alpha_{s}}$.
@@
$RS(G)\leq\frac{\sum_{s\in S,s>1}\frac{\pi_{s}}{\alpha_{s}}}{T(\sigma)}\leq\frac{1}{\alpha_{2}}\frac{\sum_{s\in S,s>1}\pi_{s}}{T(\pi)}\leq\frac{1}{\alpha_{2}}.$
@@
First, we show that $G$ is a secure restaking graph. Note that the large validator, indexed with $m^{2}+1$} and with stake $\sigma_{m^{2}+1}=2m$ does not want to participate in any attack, as she is losing at least the total value available across all services. No service $s\in\{1,2,...,m\}$ can be attacked by all validators securing it other than the large validator, since the service security parameter is $\alpha_{s}=\frac{1}{2m+1}$, while all small validators indexed ((s1)m+1,(s1)k+2,,sm)$($s-1)m+1,(s-1)k+2,...,sm) make up stake $1$ in total. The total available stakes securing a service $s$ is $2m+1$.
@@
  • and for each service we need to add extra stake at least $m$.
@@
It is easy to show that $O(m^{2})$ total stake is enough to add to validators in $V$, so that the resulting stake vector $\sigma\in\mathcal{W}(G)$. It can, for example, be done by adding the stake $2m^{2}-m$ to the validator $m^{2}+1$ so that it total stake become $(2m+1)m$. Then splitting this stake into $m$ stakes of $2m+1$ each in each of $m$ PoS protocols. Small stakers in these protocols can not attack any of the services even if they all coordinate, while the large staker does not find any attack profitable. This proves the claim of the theorem, $RS(G)\in\Theta(m)$.

### v3

@@ Abstract
We compare the total capital efficiency of secure restaking and Proof-of-Stake (PoS) protocols. First, we consider the sufficient condition for the restaking graph to be secure. The condition implies that it is always possible to transform such a restaking graph into separate secure PoS protocols. Next, we derive two main results: upper and lower bounds on the required extra stakes to add to the validators of the secure restaking graph to be able to transform it into secure PoS protocols. In particular, we show that the restaking savings compared to PoS protocols can be very large and can asymptotically grow as a square root of the number of validators. We also study a complementary question of aggregating secure PoS protocols into a secure restaking graph and provide lower and upper bounds on the PoS savings.
@@ 1 Introduction
The cryptoeconomic security of blockchains has evolved from energy-intensive Proof-of-Work to more efficient Proof-of-Stake (PoS), and most recently to restaking protocols  [^1] Total value locked in all restaking protocols amounts to more than $\mathdollar 20B$https://defillama.com/protocols/restaking, accessed on January, 2026.. Restaking allows validators to reuse their staked capital across multiple services or projects, creating a bipartite graph between validators and services. This reuse promises significant capital efficiency gains over running independent PoS protocols for each service – but it also introduces complex interdependencies and potential security risks.
@@
It can be argued that restaking gives a boost to newer projects in attracting capital for their cryptoeconomic security. Cryptoeconomic security refers to a security derived from staked parties being slashed their stakes in case of misbehavior. Restaking seems to be attractive for the validators of the original protocol as well. In particular, restaking allows them to reuse idle computational power and storage allocation for additional rewards from other projects. The original protocol, for which the validators are staked, may have requirements that do not exhaust all the resources of the validators. Setting low requirements is usually done for the purpose of decentralization [^2] Check out Ethereum network plans to resolve this asymmetry between validator resources using rainbow staking https://ethresear.ch/t/unbundling-staking-towards-rainbow-staking/18683.. Another potential reason for such availability might be economies of scale for establishing further validation services.
@@
The simplest way to verify the security of the restaking uses a trivial cost-benefit analysis. In this analysis, the total amount of lost stakes of attackers is compared to the total value obtained from the attack of a set of validators 333 [3] discusses ways to estimate the value gained from attacking a protocol and cryptoeconomic security in more detail.. The same cost-benefit analysis can be used to check the security of any PoS protocol, in which a single project/service is secured by multiple validators. In fact, the cost-benefit analysis for restaking is a direct generalization of the cost-benefit analysis for PoS protocols. This allows a comparison of restaking and multiple PoS protocols, which we develop in this paper.
@@
Conceptually, the PoS protocol is simpler and a more natural concept than restaking. Hence, we are interested in checking comparative benefits of restaking protocols. One particular attractive point of restaking protocols can be the lower total staking requirements than that of PoS protocols, since there are opportunity costs associated with locking up stake for security reasons. A key open question is whether restaking truly reduces total staking requirements compared to equivalent secure PoS protocols, and if so, by how much.
@@
Most PoS protocols pay for the opportunity cost of locking up a capital by rewarding stakers with newly issued tokens, which leads to inflation. Hence, minimizing the total stake amount while maintaining security can be a reasonable goal in both restaking or PoS protocols. Starting with a secure restaking protocol and initial stake endowments for all validators, we measure how much extra stake would be required if validators decided to form separate Proof-of-Stake protocols with their stake amounts, dividing them across the same set of projects they secure in the original restaking graph.
@@
The security parameter of a service establishes what share of the total stake securing the service is needed to successfully attack this service. Typical values of security parameters are from the set $\{\frac{1}{2},\frac{1}{3},\frac{1}{4}\}$  [^4] These values are mostly coming from the Byzantine Fault Tolerant protocols for achieving the consensus over the state. However, the values can be lower too..
Let us first check that the restaking graph is secure. It is easy to see:
@@
  • the second validator alone can not attack any service: it is not securing the first service, while the second service has large enough security parameter.
  • finally, both validators can attack both services but they lose the same amount of stake as they gain the value.
@@
Figure 1: Example of a secure restaking graph.
We only allow all validator stakes to (weakly) increase. This captures a real-life consideration in which validators’ stakes can not be taken away, but their stakes can be increased through financing when needed. In a sense, we look into over-collateralization of the whole network needed to allow separate secure PoS protocols. Similarly, [4] investigates the over-collateralization of individual validators (instead of the full set of validators as we do in this paper) that allows robustness of the restaking network security. Most of our results can be extended to the model with individual over-collaterization, obtaining the same bounds.
If both increasing and decreasing of validator stakes were allowed, it would be trivial to construct a restaking graph with minimum total stake. This amount is equal to the sum of all project values, and secure PoS protocols for all projects can be constructed by dividing the stakes of the validators. See Proposition 2 and discussion before it.
@@
    We check the sufficient condition for a secure restaking graph, specified by the EigenLayer project [9]. This sufficient condition is the only currently available (generic) condition for secure restaking graph, that can be checked in polynomial time in the size of the restaking graph. Restaking graphs that satisfy the sufficient condition do not save on the total required stake. In particular, it is possible to divide the initial stakes of the validators across projects they validate, to obtain secure PoS protocols.
@@
    To further extend the analysis, we define restaking savings of a fixed secure restaking graph. It is equal to the extra stake required, relative to the original total available stake in the graph, to be able to obtain secure PoS protocols through the division of stake described above.
@@
    We study a complementary question of aggregating given secure PoS protocols into a secure restaking graph. When aggregating, each validator adds stakes across projects that they secure in separate PoS protocols, resulting in a restaking graph. Similarly to the restaking savings, we define PoS savings as the total additional stake required to add validators in the aggregate restaking graph to make it secure. We first construct an example showing that the resulting restaking graph is not secure, in fact, at least an additional stake equal to the original total stake is required to make it secure, thus showing a lower bound of $1$ on PoS savings. We also provide an upper bound on PoS savings as a function of the underlying restaking graph, which matches a lower bound example for a special set of graphs.
@@
[7] introduced the first cryptocurrency Bitcoin and the concept of PoW, in which the random sampling of the next block proposer is done proportionally to a hashrate of a miner, by solving a computational puzzle. [8] provides a detailed discussion and economic comparison on PoW and PoS concepts. [3] discusses ways to estimate the value gained from attacking a protocol and cryptoeconomic security in more detail. The idea of restaking was originally proposed in the whitepaper of the Eigenlayer project, [9]. A similar concept exists in traditional finance literature under the name of rehypothecation[5]. The interdependencies between protocols and validators introduced by restaking create new risks, which have been modeled and analyzed in the original Eigenlayer whitepaper, as well as in a follow up paper [4], in which the authors study robust restaking networks, resistant to a failure of a fraction of validating stakes and cascading effects. The authors consider over-collaterization factor of individual validators. They investigate out how large should this uniform over-collaterization factor be so that the originally secure restaking graph stays robust under these crashes. Our over-collaterization over the total stake amount can also be reduced to individual over-collaterization factors, as we argue after the Theorem 2.
[2] studies a setting in which the value of attacked set of services is not the sum of values in the set, but a concave function. [10] studies elastic restaking services, in which validators allocate stakes to services that may make up more than their initial stake. This allows for a more expressive model than the original restaking model where the whole stake of each validator is locked up for all services. The work also considers a typical distributed computing setting in which only a fraction of validators is Byzantine. [6] studies risks and benefits of combining different restaking services. [1] studies economic security of different protocols as a function of streaming rewards.
@@ 2 Notation and Model
Our notation follows closely [4]. A restaking graph $G=(S,\pi,\alpha,V,\sigma,E)$ consists of the following:
  • $S$ denotes the set of $m$ services $\{1,...,m\}$ (we will refer projects as services from now on throughout the paper),
  • $\pi\in\mathbb{R}^{m}_{+}$ denotes values of services,
@@
Let $N_{G}(s)$ denote the neighborhood of the service $s\in S$ in the restaking graph $G$, i.e. $N_{G}(s):=\{v\in V:(s,v)\in E\}$, is the set of validators staked on, and, therefore, securing a service $s$. Similarly, $N_{G}(v)$ denotes the neighborhood of the validator $v\in V$, i.e., $N_{G}(v):=\{s\in S:(s,v)\in E\}$, is the set of services validator $v$ is staked on.
@@
A division of stake $\sigma_{v}$ of validator $v$ across services it secures in $G$ induces a vector $c^{v}\in\mathbb{R}^{|N_{G}(v)|}_{+}$, such that, $\sum_{s\in N_{G}(v)}c^{v}_{s}=\sigma_{v}$. We say that division of stake is secure if in the corresponding Proof-of-Stake protocols, that represent star graphs centered at the service, we have guaranteed security. That is, for each service $s$ with security parameter $\alpha_{s}$, a vector of stakes $\{k\in N_{G}(s):c^{k}_{s}\}$ is secure, which means that there is no profitable attack by validators having stakes $\{k\in N_{G}(s):c^{k}_{s}\}$. Let $\mathcal{W}(G)$ denote all initial stake vectors $\sigma$ for which such division is possible. To simplify notation, here we implicitly assume that the stake vector is not part of a restaking graph $G$, and only keep its combinatorial structure together with the $\pi$ and $\alpha$ vectors.
@@ 3 Results
In this section, we first check the sufficient condition (1) and its implications. The implications motivate to define restaking and Proof-of-Stake savings. Last, we obtain lower- and upper-bound results for them both types of savings.
@@ 3.1 Restaking Savings
stake to a service $s$ it secures in $G$, in the PoS protocol. Such allocation is possible because of (1). Each service $s$ is assigned a total stake amount of at least $\frac{\pi_{s}}{\alpha_{s}}$, since
@@
Note that the value $\frac{\pi_{s}}{\alpha_{s}}$ is the minimum required total stake in a PoS protocol for a service $s$ with value $\pi_{s}$ and security parameter $\alpha_{s}$ that guarantees that there is no profitable attack, independent of how this total stake is divided between validators. For any smaller amount $t<\frac{\pi_{s}}{\alpha_{s}}$, there exists a distribution of the total stake $t$ among any number of validators such that there is a profitable attack.
On the other hand, if we consider suitable division of stakes between validators, we obtain a much weaker requirement to make secure PoS protocols. Namely, we obtain that there is a minimum total stake requirement to ensure that there is a division in secure PoS protocols.
@@
Another consideration is that the restaking graph with the minimum total stake in $\mathcal{W}(G)$ is not necessarily secure, while we will focus on secure restaking graphs from now on.
We are ready to define the main metric through which we measure capital efficiency of a secure restaking graph, restaking savings:
@@
where indicator function $\mathbb{I}(P)=1$ whenever the statement $P$ is true.
@@
    the vector of stakes $(c_{1},...,c_{d_{G}(s)})$ is equal to $(0,...,0,\pi_{s}+\sigma_{v},0,...0)$ for any $v$. This construction is secure for a similar reason as above.
@@
Next, we show $RS(G)\leq K$. Consider a cover of vertices in $V$ by neighborhoods of services $s\in R\subseteq S$ that has the minimum maximum incidence number $K$. For each neighborhood $N_{G}(s)$, where $s\in R$, use construction 3: original $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ stakes from the restaking graph to secure service $s$ in the PoS protocol. To be able to do that, we need to add at most K1K-1 times all $\sigma_{v}$ stakes, for any $v\in V$. In this way, we utilize all $\sigma_{v}$ stakes and secure services in the cover set $R$. For service $s$ in $S\setminus R$, we add $\pi_{s}$ stake once to any of its neighbors in $N_{G}(s)$ and use construction $(0,...,0,\pi_{s},0,...,0)$ to secure service $s$ in a corresponding PoS protocol. This makes sure we have secured all services and utilized all stakes of all validators. Then, $RS(G)\leq(K-1)+1=K$. ∎
Note that $RS(G)\leq K$ in particular implies that $RS(G)\leq\max_{v\in V}(d_{G}(v))$, as $K$ is upper bounded by $\max_{v\in V}(d_{G}(v))$. There are examples where $K=\max_{s\in S}(d_{G}(s))$, implying that the upper bound can be as high as $n$. However, next, we show an upper bound as a function of the number of validators, which is asymptotically much lower than $n$.
@@
The proof relies on a greedy covering of high-degree services. The proof combines constructions from the proof of Proposition 3. Initialize the set $S_{L}:=\emptyset$ and $V_{L}:=\emptyset$. The following procedure is repeated until it is no longer possible: At each step, find a service $s$ in $S\setminus S_{L}$ such that its degree in the set $V\setminus V_{L}$ is at least $\sqrt{n}$. Update the set $S_{L}:=S_{L}\cup s$ and $V_{L}:=V_{L}\cup N_{G}(s)$. Note that after at most $\sqrt{n}$ steps, it is not possible to find a new service $s$ and, hence, the procedure stops. That is, for the resulting set of services $S_{L}$, we have $|S_{L}|\leq\sqrt{n}$. This, in particular, implies that the maximum outdegree of a validator $v$ in $V_{L}$ to $S_{L}$, denoted by $d_{S_{L}}(v)$, is upper bounded by $\sqrt{n}$.
@@
Note that in the proofs of Proposition 3 and Theorem 1, we did not use any property of $\alpha$ values. Both upper bounds use combinatorial features of the underlying restaking graph $G$ and hold for any values of security parameters $\alpha$. Next, we show an upper bound that is a function of the security parameters $\alpha_{s}$. We show the following upper bound on $RS(G)$:
@@
Assume that services are labeled in the way that the security parameters are sorted in increasing order $\alpha_{1}\leq\alpha_{2}\leq...\leq\alpha_{m}$. Then, $RS(G)\leq\frac{1}{\alpha_{2}}$.
@@
For the service $s=1$, with the security parameter $\alpha_{1}$, allocate the original stakes of the restaking graph in the PoS protocol: $(c_{1},...,c_{d_{G}(1)})=(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(1)}})$, where $N_{G}(1)=\{v_{1},...,v_{d_{G}(1)}\}$ is the set of validators staked on service $s=1$. This PoS protocol is secure by the assumption tjat tje original restaking graph is secure. Repeat the following procedure for all services $s>1$:
@@
    use construction $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ to secure service $s$ in PoS protocol, where $N_{G}(s)=\{v_{1},...,v_{d_{G}(s)}\}$ is the set of validators staked on service $s$,
@@
    set all stake values in the vector $(\sigma_{v_{1}},...,\sigma_{v_{d_{G}(s)}})$ to 0.
Each PoS derived at step 2. is secure, as the total stake allocated for a service $s>1$ is at least $\frac{\pi_{s}}{\alpha_{s}}$. This is an observation made in the proof of Proposition 1.
@@
$RS(G)\leq\frac{\sum_{s\in S,s>1}\frac{\pi_{s}}{\alpha_{s}}}{T(\sigma)}\leq\frac{1}{\alpha_{2}}\frac{\sum_{s\in S,s>1}\pi_{s}}{T(\pi)}\leq\frac{1}{\alpha_{2}},$
finishing the proof of the proposition claim. ∎
@@
2221/31/31/31/31/31/31/31/31/36 $S$ $V$
Figure 2: Example construction for $m=3$.
First, we show that $G$ is a secure restaking graph. Note that the large validator, indexed with $m^{2}+1$} and with stake $\sigma_{m^{2}+1}=2m$ does not want to participate in any attack, as it is losing at least the total value available across all services. No service $s\in\{1,2,...,m\}$ can be attacked by all validators securing it other than the large validator, since the service security parameter is $\alpha_{s}=\frac{1}{2m+1}$, while all small validators indexed ((s1)m+1,(s1)k+2,,sm)$($s-1)m+1,(s-1)k+2,...,sm) make up stake $1$ in total. The total available stakes securing a service $s$ is $2m+1$.
@@
  • for each service we need to add extra stake at least $m$.
@@
It is easy to show that $O(m^{2})$ total stake is enough to add to validators in $V$, so that the resulting stake vector $\sigma\in\mathcal{W}(G)$. It can, for example, be done by adding the stake $2m^{2}-m$ to the validator $m^{2}+1$ so that its total stake becomes $(2m+1)m$. Then, assume the validator splits this total stake into $m$ equal stakes of $2m+1$, for each $m$ PoS protocols. The small stakers in these protocols can not attack any of the services even if they all coordinate, while the large staker does not find any attack profitable. This proves the claim of the theorem, $RS(G)\in\Theta(m)$.
@@
From the proof of Theorem 2 we can also observe that at least one of the individual validator stakes should be increased by a factor of $\Theta(m)$. Therefore, if restaking savings were to be measured like over-collaterization factor of individual validators, as in [4], the asymptotics would be the same. This adds to the strength of the lower bound example.